### v1

# Foundation Models for the Brain and Body

Abstract
Subject-specific distribution shifts represent an important obstacle to the development of foundation models for EEG decoding. To address this, we propose Subject-Conditioned Layer, an adaptive layer designed as a drop-in replacement for standard linear or convolutional layers in any neural network architecture. Our layer captures subject-specific variability by decomposing its weights into a shared, subject-invariant component and a lightweight, low-rank correction unique to each subject. This explicit separation of general knowledge from personalized adaptation allows existing models to become robust to subject shifts. Empirically, models equipped with our layer outperform both a shared-weight-only model (subject-agnostic model) and the average of individually trained subject-specific models. Consequently, the Subject-Conditioned Layer offers a practical and scalable path towards building effective cross-subject foundation models for EEG.
1 Introduction
Developing a foundation model for decoding complex brain signals like Electroencephalography (EEG) is a central challenge in neuroscience and machine learning [2]. Progress is slowed by the inherent properties of EEG data: low signal-to-noise ratio, non-stationarity, and, most critically, high inter-subject variability [35]. Neural patterns vary so widely across individuals, and even within the same individual over time, that data is effectively fragmented, exacerbating the problem of data scarcity [3, 25]. Consequently, large models trained on pooled data from many subjects often fail to outperform simpler models tailored to individuals, stalling progress towards generalizable solutions.
This situation contrasts sharply with fields like computer vision (CV) and natural language processing (NLP). While these fields have also relied on domain-specific architectures, the Transformer [34] provided a universal paradigm that unlocked predictable scaling laws, enabling massive foundation models to emerge [13, 37]. EEG decoding has yet to benefit from similar scaling. The challenge of cross-subject generalization has kept effective dataset sizes small, preventing the successful application of large-scale, general-purpose architectures. As a result, the field remains dominated by specialized models that have not yet converged to a standard, foundational approach [2, 12].
To bridge this gap, we address the problem of subject-specific distribution shifts, which we identify as the primary obstacle to building foundation models for multi-subject problems such as EEG classification. We propose Subject-Conditioned Layer, an approach that decomposes the parameters of a neural network into two components: a set of base weights shared across all subjects and a subject-specific, low-rank correction that efficiently captures individual variability. The Subject-Conditioned Layer module is designed as a drop-in replacement for standard linear or convolutional layers, making it universally applicable to a wide range of model architectures with applications in EEG decoding and beyond.
Our contributions are threefold: We introduce Subject-Conditioned Layer, a novel adaptive layer that explicitly separates shared knowledge from subject-specific signatures and is designed as a drop-in replacement for standard linear and convolutional layers in any neural network architecture. We provide qualitative evidence through embedding visualization that our layer successfully disentangles subject-invariant neural patterns from individual subject signatures, confirming the mechanism behind our method’s success. Through a comprehensive evaluation on the BCI Competition IV datasets, we demonstrate that both CNN and ViT like architectures equipped with Subject-Conditioned Layer outperform strong baselines: (i) a subject-agnostic model, (ii) the average of subject-specific models, and (iii) the average of subject-specific LoRA models. The code is publicly available  [^1] https://github.com/timonkl/SubjectConditionedLayer.
Figure 1: Latent Representations Colored by Subject ID vs. Target Labels. The top row compares t-SNE projections of representations derived from (Left) the shared weights ($W_{\text{general}}$), (Middle) the subject-specific weights ($W_{s}$), and (Right) the fused representations from the complete Subject-Conditioned Layer. In this row, all points are colored by subject ID. The bottom row presents the same data points, but colored according to their ground-truth target labels. This comparison clearly shows that the sub-clusters formed by our full model (top right) correspond directly to the target classes (bottom right), demonstrating that the model learns discriminative, task-relevant features within each subject’s representation space. All representations are obtained from the classification token predicted by the Patched Brain Transformer, with a linear classification head used for visualization clarity.
2 Related Work
Our proposed method addresses the challenge of inter-subject variability in EEG decoding by learning both shared and subject-specific representations. This approach is closely related to established paradigms nonlinear mixed-effects (NLME) modeling and in machine learning, primarily Multi-Task Learning (MTL) and Mixture of Experts (MoE).
Nonlinear Mixed-Effects Modeling. Our approach can be viewed through the lens of nonlinear mixed-effects (NLME) models [21, 22]. In this framing, the shared network $W_{\text{general}}$ learns the fixed, population-level effects, while the subject-specific low-rank adapters $W_{s}$ capture the random effects, parsimonious deviations for each individual. The low-rank constraint on the adapters is analogous to the low-dimensional random effects assumption in classical NLME, which encourages partial pooling of information and borrows statistical strength across subjects. This perspective provides a principled justification for using LoRA [14] as a parameter-efficient method to model subject-specific variations, which is crucial when per-subject data is limited.
Multi-Task Learning. In MTL, a model is trained on several related tasks simultaneously, with the goal of improving performance by leveraging shared information across tasks [5, 23, 9]. Viewing the decoding for each subject as a distinct task, our model fits directly into the MTL framework. It employs a form of hard parameter sharing, where the general weight matrix, $W_{\text{general}}$, learns a common feature space shared across all subjects. The subject-specific weights, $W_{s}$, act as task-specific layers that adapt this shared representation to the unique neural patterns of each individual. This architecture allows the model to learn robust, generalizable features from the entire population while still retaining the flexibility to specialize for each subject [38].
Mixture of Experts. MoE models utilize multiple expert networks and a gating mechanism that determines which expert processes a given input [15]. Sparsely-gated MoE models, in particular, activate a subset of experts for each input, enabling the scaling of model capacity without a proportional increase in computation [31, 10]. Our subject-aware layer can be interpreted as a specialized, non-learned MoE. Here, the subject identity serves as the gating signal, deterministically routing the input to its corresponding expert, the pathway defined by $W_{s}$. Unlike traditional MoE, our gating is not a learned function but is provided by the input via the mask M. This simplifies the model by removing the need to learn a routing policy and directly addresses the known source of variation (i.e., subject identity).
Dynamical low-rank. Since our method explicitly models subject-specific variability as lightweight low-rank corrections, it naturally connects to the broader line of dynamical low-rank research [26, 28, 29]. An exciting direction for future work is to explore whether subject adaptation in EEG can benefit from recent advances in geometry-aware training [36, 30], and robust compression strategies [29], thereby unifying personalization with efficient large-scale deployment.
3 On the Difficulty of Multi-Subject Training
We consider a classification problem where data is collected from multiple subjects. Since the signal contains subject-specific characteristics, our goal is to learn a general model that captures class-discriminative structures across subjects. Let $\mathcal{X}_{s}=\{x_{sj}\}_{j=1}^{M_{s}}$ and $\mathcal{Y}_{s}=\{y_{sj}\}_{j=1}^{M_{s}}$ contain subject-specific features and labels, respectively, for subjects $s\in\{1,\dots,N\}$. In short, we write $\mathcal{D}=(\mathcal{X},\mathcal{Y})=\bigcup_{s=1}^{N}\mathcal{D}_{s}=\bigcup_{s=1}^{N}(\mathcal{X}_{s},\mathcal{Y}_{s})$.
A single linear layer trained for all subjects implicitly estimates $p(\mathcal{Y}|\mathcal{X})$, but cannot properly capture subject-specific features because it lacks a mechanism to adapt its weights per subject. It can only represent the shared average structure. We address this with a probabilistic framework that models subject-specific effects. Our main goal is to learn the subject-independent classifier $p(\mathcal{Y}|\mathcal{X})$ that for a given $\mathcal{D}$ marginalizes over $s$,
$p(\mathcal{Y}\mid\mathcal{X})=\textstyle\sum_{s=1}^{N}p(\mathcal{Y}\mid\mathcal{X},s).$ (1)
Further, assume that we have latent random variables $\mathcal{Z}=\xi(\mathcal{X},s)$ as intermediates on the map from $\mathcal{X}$ to $\mathcal{Y}$, where we define $\xi$ as a transformation to latent representation. Then we can write
$p(\mathcal{Y}\mid\mathcal{X},s)=p(\mathcal{Y}\mid\mathcal{Z},s)\;p(\mathcal{Z}\mid\mathcal{X},s).$ (2)
Concerning the first term of Equation (2) we assume that
$p(\mathcal{Y}\mid\mathcal{Z},s)=p(\mathcal{Y}\mid\mathcal{Z}),$
which states that once the latent representation $\mathcal{Z}$ is known, the participant index $s$ and the datapoint $x_{sj}$ provide no additional information about the label $y_{sj}$. This conditional independence assumption implies that $\mathcal{Z}$ serves as a sufficient statistic of $x_{sj}$ to predict $y_{sj}$. A key step is to verify whether this property holds for the proposed method.
The second term of Equation (2), $p(\mathcal{Z}\mid\mathcal{X},s)$, describes how the latent representation is learned. Specifically,
$\xi(\mathcal{X},s)\sim p(\mathcal{Z}\mid\mathcal{X},s),$ (3)
i.e., $\xi$ is treated as a stochastic latent variable in the representation space $\mathcal{Z}$, sampled conditionally on the input $\mathcal{X}$ and the participant index $s$. The subject-specific map $\xi(\mathcal{X},s)$ is parameterized by a set of general weights and a subject-specific low-rank correction.
4 Method
Figure 2: Subject-Conditioned Layer. The input $X$ is processed by a shared general weight matrix $W_{\text{general}}$. For each subject, a subject-specific low-rank adapter, parameterized by matrices $A_{s}$ and $B_{s}$ of rank $r$, computes a correction that is added to the output of the general weight matrix, yielding the layer output $\bar{X}$.
4.1 Subject-Conditioned Layer
We propose a novel subject-aware layer for multi-subject machine learning tasks as
$\bar{X}=\sigma\left(XW_{\text{general}}^{\top}+\textstyle\sum_{s=1}^{N}(M_{s}\cdot X)W_{s}^{\top}\right),$ (4)
where $W_{\text{general}}\in\mathbb{R}^{m\times n}$ and $W_{s}\in\mathbb{R}^{m\times n}$ are shared and subject-specific weight matrices, respectively. The mask matrices $M_{s}$ with binary entries extract subject-$s$-specific samples from the batch data $X$.
Note that this approach may lead to a reduction in batch sizes for subject-dependent weights $W_{s}$, if the number of subjects $N$ increases.
As a remedy, adapting the total batch size was proposed in [31].
4.2 Low Rank approximation for Subject-Dependent Weights
Low-rank adaptation (LoRA) [14] is a well-established method for fine-tuning models in few-shot settings. Compared to full-rank training, LoRA shows superior performance on limited data [27], while also substantially reducing the number of training parameters and the associated memory consumption. We define subject-specific weights as low-rank matrix decompositions
$W_{s}^{\top}:=A_{s}B_{s},$ (5)
with $A_{s}\in\mathbb{R}^{n\times r},\;B_{s}\in\mathbb{R}^{r\times m}$ and rank $r$. Similarly to [14], we initialize $A_{s}=\mathcal{N}(0,\sigma^{2})$, and $B_{s}=0$ for all subjects $s$.
4.3 Regularization of Subject-Dependent Weights
A central challenge is to ensure that the general weight matrix $W_{\text{general}}$ learns shared patterns while the subject-specific weights $W_{s}$ only capture individual deviations. To prevent $W_{s}$ from learning general representations, we employ a regularization strategy. First, we constrain the complexity of the subject-specific corrections by enforcing a relatively small rank $r$, as defined in Equation 5. See Tables 2 and 3 for exact values.
Second, we introduce an optimization-based regularization method by scaling $W_{s}$ with $\frac{\alpha}{r}$, where $\alpha$ is a hyperparameter. This scaling adjusts the effective learning rate of the adapter matrices relative to the general weights. For $\alpha<r$, the adapters are updated conservatively, encouraging reliance on $W_{\text{general}}$. In contrast to $\alpha>r$, where the adapters are emphasized.
The weights $W_{s}$ could also be regularized by adding a weight regularization term (e.g., an $L_{2}$ penalty) to the loss function. We did not pursue this here, opting instead to demonstrate the capabilities of the LoRA only approach.
5 Extension to Convolutional Layers
While the formulation in Equation 4 is presented for linear transformations, it can be seamlessly extended to convolutional layers. A convolution operation is itself a linear transformation, where the weight matrix $W$ is replaced by a convolutional kernel $W$. The Subject-Conditioned convolutional layer can thus be expressed as:
$\bar{X}=\sigma\left(X*\mathcal{W}_{\text{general}}+\textstyle\sum_{s=1}^{N}M_{s}\cdot(X*\mathcal{W}_{s})\right),$ (6)
where * denotes the convolution operation.
To implement the low-rank adaptation (Section 4.2) for the subject-specific kernels, we decompose the operation $X*W_{s}$ into two sequential convolutions, effectively factorizing the kernel itself. This is analogous to the matrix decomposition in Equation 5. Specifically, the subject-specific update is computed by:
$\mathcal{W}_{s}:=A_{s}B_{s},$ (7)
where $A_{s}$ and $B_{s}$ are the low-rank convolutional kernels. The first kernel, $A_{s}$, reduces the channel dimension from $C_{\text{in}}$ to the rank $r$ while preserving the spatial kernel size. The second kernel, $B_{s}$, is a $1\times 1$ convolution that projects the features from the rank $r$ space back to the $C_{\text{out}}$ channel dimension. Initialization is equal to the linear case, see Section 4.1. This convolutional decomposition allows for parameter-efficient adaptation for each subject while maintaining the inductive biases of convolutional layers.
6 Data
We conduct our experiment on two Brain Computer Interface datasets from BCI Competition IV [32]. The data are publicly available [^2] http://bnci-horizon-2020.eu/database/data-sets with the MOABB-Package [7, 1]. The BCI Competition IV 2a dataset [4] consists of 22-channel EEG recordings from nine subjects. For each subject, data were collected in two sessions (training and evaluation), with 288 trials per session. The labels are balanced across four motor imagery classes: left hand, right hand, foot, and tongue. The BCI Competition IV 2b dataset [20, 19] consists of 3-channel EEG recordings from nine subjects. For each subject, data were collected across five sessions, three for training and two for evaluation, totaling between 680 and 760 trials. The labels are balanced between the classes: left hand and right hand.
7 Results
We evaluate our proposed method, Subject-Conditioned Layer, by integrating it into two distinct EEG decoding architectures: EEGNeX [6], a compact Convolutional Neural Network (CNN), and the Patched Brain Transformer (PBT) [17], a Vision Transformer (ViT) based architecture [8] that reflect recent trends in the field [18, 16]. These models demonstrate the versatility of Subject-Conditioned Layer across diverse architecture types. All experiments were conducted on the BCI Competition IV-2a and IV-2b datasets. Further details on the experimental setup are in Appendix A.
7.1 Qualitative Analysis
To confirm our hypothesis that our Subject-Conditioned Layer learns to decouple shared from subject-specific information, we visualize the PBT’s latent representations using t-SNE [33], see Figure 1. The visualizations reveal a clear separation of the learned representations. As shown in the figure, embeddings from the shared weights ($W_{\text{general}}$) are unstructured by subject (left), whereas those from the subject-specific weights ($W_{s}$) form distinct, individual-specific clusters (middle). The full model (right) integrates both aspects: it preserves the subject-level clustering while also forming meaningful sub-clusters corresponding to target labels, Figure 1, demonstrating that it retains task-relevant information. This demonstrates that the low-rank corrections $W_{s}$ have effectively isolated the idiosyncratic signatures of each subject. Together, these visualizations provide compelling evidence for our method’s proposed mechanism.
7.2 Numerical Results
To validate our approach, we compare Subject-Conditioned Layer against several modeling strategies, with results summarized in Table 1. All numerical results are reported as the mean and standard deviation of three reproducible runs with random seeds $\{1,2,3\}$.
First, to establish the necessity of subject-specific adaptation, we compare against a Subject-Agnostic Model. This baseline is a standard architecture trained on data collected from all subjects, without personalization. Our method surpasses this baseline, confirming that a one-size-fits-all approach is insufficient for high performance in EEG decoding, Table 1.
Next, we demonstrate the benefit of knowledge sharing by comparing our method against the common practice of training fully independent Subject-Specific Models. This baseline trains a separate model from scratch for each individual. In Table 1, we report the mean performance across all subjects (the complete results are provided in Table 4). Subject-Conditioned Layer consistently outperforms the average of those independently trained models. This finding highlights that the shared backbone ($W_{\text{general}}$) enables effective knowledge transfer across subjects, yielding better generalization than isolated training on limited single-subject datasets.
Finally, to validate our specific architectural design, we introduce a crucial parameter control: Subject-Specific LoRA Models. For each subject, a separate low-rank model is trained from scratch. This baseline allows us to test if the performance gain is simply an artifact of having a certain number of trainable parameters per subject. Subject-Conditioned Layer’s superiority in this comparison is a critical finding. It proves that our hybrid architecture is more effective than training isolated, parameter-efficient models. This validates that the structure of our approach is important for its performance.
Table 1: Numerical Results. Performance comparison across model variants: a subject-agnostic model (no personalization), subject-specific models (trained independently per subject), subject-specific LoRA models (entirely low-rank per subject), and Subject-Conditioned Layer.
Method Total Active BCI Comp. IV2a BCI Comp. IV2b
# Params # Params (4 Classes) (2 Classes)
EEGNeX Subject-Agnostic [6] 155 972 55 972 55.00% $\pm$ 0.66% 74.28% $\pm$ 0.39%
EEGNeX Subject-Specific 155 972 55 972 59.36% $\pm$ 14.42% 75.92% $\pm$ 14.32%
EEGNex Subject-Specific LoRA 181 124 10 980 53.60% $\pm$ 2.91% 69.64% $\pm$ 0.27%
EEGNeX Subject-Conditioned Layer 134 884 64 740 64.72% $\pm$ 0.56% 76.48% $\pm$ 0.25%
PBT Subject-Agnostic [17] 1 867 460 867 460 50.82% $\pm$ 0.85% 75.49% $\pm$ 1.10%
PBT Subject-Specific 1 867 460 867 460 44.16% $\pm$ 12.19% 75.63% $\pm$ 14.23%
PBT Subject-Specific LoRA 1 680 736 135 712 25.81% $\pm$ 0.52% 53.70% $\pm$ 0.80%
PBT Subject-Conditioned Layer 1 480 612 935 588 54.51% $\pm$ 0.30% 76.36% $\pm$ 0.23%
8 Limitations & Outlook
Our work focuses on introducing Subject-Conditioned Layer and empirically demonstrating its effectiveness on the BCI Competition IV 2a and 2b datasets. However, our study has some limitations, which suggest directions for future research.
A limitation is the framework’s dependence on subject-specific data to train the personalized adapters, $W_{s}$. This means that for previously unseen subjects, predictions can only be made using the shared weights, $W_{\text{general}}$. Although we have shown that such personalization enhances performance, further analysis is required to fully assess the model’s generalization capabilities and its viability for practical deployment. Subject-Conditioned Layer is well-suited for a pre-training and fine-tuning paradigm. One could first train the shared weights $W_{\text{general}}$ across a large population. Subsequently, for a new subject, one could freeze the powerful shared model and efficiently fine-tune only a lightweight subject-specific adapter $W_{s}$. This iterative application, coupled with scaling, could lead to a Foundation model for EEG Decoding. To tackle the challenge of unseen subjects, future research could investigate the geometry of the learned adapter space. By comparing the adapters of different subjects (e.g., via their similarity), we could potentially discover subject clusters or archetypes. This could allow us to initialize a new subject adapter from a pre-computed group adapter, enabling faster personalization or even zero-shot prediction, thereby greatly enhancing the practical utility of our method.
Funding Acknowledgements
This project has received funding from the European Regional Development Fund (grants timingMatters and IntelAlgen) under the European Union’s Horizon Europe Research and Innovation Program, from the German Research Foundation DFG within GRK 2297 ‘Mathematical Complexity Reduction’, and from the German Federal Joint Committee (Grant 01VSF23017), as well as German Research Foundation DFG (Grant 537063406), which we gratefully acknowledge.
Appendix A Experiment Design
A.1 EEGNeX
We used the EEGNeX implementation from the Braindecode package [11, 24]. In this model, we replaced all standard convolutional layers with the convolutional version of our method, Subject-Conditioned Layer (see Appendix 5). We made an exception for the depthwise convolutional layer (DepthwiseConv2D) in Block 3 due to implementation details.
As the original hyperparameters for EEGNeX were not specified, we tuned the baseline learning rate by sweeping over the values $\{10^{-3},\;10^{-4}\}$. We want to remark that the difference between the numbers reported in Table 1 and those in the original EEGNeX paper stems from different evaluation methods. Here, we report the numbers for cross-session analysis from a 512-timepoint long interval.
The final hyperparameters are reported in Table 2:
Table 2: Hyperparameters EEGNeX.
Hyperparameter Value
Band-pass filter [8, 45]
Optimizer AdamW
Learning rate $1\times 10^{-3}$
Weight decay 0.01
AdamW betas (0.9, 0.999)
Batch size 64
Training epochs 100
Dropout rate 0.0
Activation function ELU
Rank adapters 4
LoRA $\alpha$ 1
Patched Brain Transformer
For the PBT, we used the official GitHub repository accompanying the publication and replaced all linear layers with our proposed method, Subject-Conditioned Layer. The only exception was for the visualization in Figure 1, where we kept the final linear layer to improve clarity.
We used the same hyperparameters as reported in the original PBT paper for all runs (Table 3). For our method’s specific hyperparameters, we set the rank $r=8$ and $\alpha=1$.
Table 3: Hyperparameters PBT.
Hyperparameter Value
Band-pass filter [8, 45]
Optimizer AdamW
Learning rate $3\times 10^{-4}$
Weight decay 0.01
AdamW betas (0.9, 0.95)
Batch size 128
Training epochs 1200
Warmup epochs 150
Dropout rate 0.1
Activation function GELU
Rank adapters 8
LoRA $\alpha$ 1
Data augmentation Time shifts
Appendix B Detailed Numerical Results for Subject-Specific Modeling
Table 4: Subject-Specific Numerical results for EEGNeX and PBT on individual subjects.
Subject EEGNeX PBT
BCI Comp. IV2a BCI Comp. IV2b BCI Comp. IV2a BCI Comp. IV2b
1 67.82% $\pm$ 6.25% 72.50% $\pm$ 1.84% 56.60 % $\pm$ 1.85% 71.35% $\pm$ 1.28%
2 46.53% $\pm$ 1.70% 52.86% $\pm$ 0.77% 25.23 % $\pm$ 1.31 % 53.10% $\pm$ 1.02%
3 77.08% $\pm$ 3.27% 51.56% $\pm$ 1.99% 55.90 % $\pm$ 3.07 % 50.73% $\pm$ 0.15%
4 53.01% $\pm$ 1.80% 94.27% $\pm$ 0.53% 39.00 % $\pm$ 1.15 % 95.31% $\pm$ 1.11%
5 36.69% $\pm$ 3.38% 86.25% $\pm$ 1.35% 26.97 % $\pm$ 0.59 % 86.98% $\pm$ 1.62%
6 40.97% $\pm$ 1.58% 81.77% $\pm$ 2.32% 34.72 % $\pm$ 2.79 % 77.81% $\pm$ 1.02%
7 67.94% $\pm$ 2.31% 70.94% $\pm$ 2.04% 47.80 % $\pm$ 7.45 % 78.52% $\pm$ 0.72%
8 68.29% $\pm$ 0.71% 86.77% $\pm$ 1.49% 55.79 % $\pm$ 3.38 % 86.46% $\pm$ 0.64%
9 75.93% $\pm$ 0.43% 83.85% $\pm$ 1.98% 55.44 % $\pm$ 0.59 % 83.02% $\pm$ 1.64%
Appendix C Compute Resources
The shown experiments are performed on a single NVIDIA Tesla V100 GPU using PyTorch. Code is publicly available at github.com/timonkl/SubjectConditionedLayer.

### v2

@@
-Foundation Models for the Brain and Body
+Mitigating Subject Dependency in EEG Decoding with Subject-Specific Low-Rank Adapters
 Abstract
-Subject-specific distribution shifts represent an important obstacle to the development of foundation models for EEG decoding.
-To address this, we propose Subject-Conditioned Layer, an adaptive layer designed as a drop-in replacement for standard linear or convolutional layers in any neural network architecture.
-Our layer captures subject-specific variability by decomposing its weights into a shared, subject-invariant component and a lightweight, low-rank correction unique to each subject.
-This explicit separation of general knowledge from personalized adaptation allows existing models to become robust to subject shifts.
-Empirically, models equipped with our layer outperform both a shared-weight-only model (subject-agnostic model) and the average of individually trained subject-specific models.
-Consequently, the Subject-Conditioned Layer offers a practical and scalable path towards building effective cross-subject foundation models for EEG.
+Subject-specific distribution shifts represent a fundamental obstacle to developing foundation models for brain decoding.
+We propose the Subject-Specific Low-Rank Adapter (SuLoRA), a drop-in replacement for standard linear or convolutional layers that captures inter-subject variability by decomposing weights into a shared, subject-invariant component and a lightweight, low-rank correction unique to each subject.
+This explicit separation enables existing architectures to become robust to subject shifts without architectural redesign.
+We evaluate SuLoRA on MEG speech perception and EEG motor imagery tasks across CNN and transformer architectures.
+In the speech decoding task, SuLoRA exceeds the baseline performance with half of the parameters.
+On motor imagery dataset, SuLoRA outperforms both subject-agnostic models and independently trained subject-specific models.
+SuLoRA offers a practical path towards effective cross-subject foundation models for brain signal applications.
+Machine Learning, ICML
 1 Introduction
-Developing a foundation model for decoding complex brain signals like Electroencephalography (EEG) is a central challenge in neuroscience and machine learning [2].
-Progress is slowed by the inherent properties of EEG data: low signal-to-noise ratio, non-stationarity, and, most critically, high inter-subject variability [35].
-Neural patterns vary so widely across individuals, and even within the same individual over time, that data is effectively fragmented, exacerbating the problem of data scarcity [3, 25].
-Consequently, large models trained on pooled data from many subjects often fail to outperform simpler models tailored to individuals, stalling progress towards generalizable solutions.
-This situation contrasts sharply with fields like computer vision (CV) and natural language processing (NLP).
-While these fields have also relied on domain-specific architectures, the Transformer [34] provided a universal paradigm that unlocked predictable scaling laws, enabling massive foundation models to emerge [13, 37].
-EEG decoding has yet to benefit from similar scaling.
+Developing a foundation model for decoding complex brain signals such as Electroencephalography (EEG) and magnetoencephalography (MEG) is a central challenge in neuroscience and machine learning (Aristimunha et al., 2025c).
+Progress is slowed by the inherent properties of EEG data: low signal-to-noise ratio, non-stationarity, and, most critically, high inter-subject variability (Wei et al., 2022).
+Neural patterns vary so widely between individuals and even within the same individual over time.
+The data are effectively fragmented, amplifying the problem of data scarcity (Banville et al., 2025; Bomatter and Gouk, 2025; Schmid et al., 2024).
+Consequently, large models trained on pooled data from many subjects often fail to outperform simpler models tailored to individuals, slowing progress towards generalizable solutions.
+This situation contrasts sharply with fields of natural language processing or computer vision.
+Although these fields have also relied on domain-specific architectures, the Transformer (Vaswani et al., 2023) provided a universal paradigm that unlocked predictable scaling laws, allowing massive foundation models to emerge (Hoffmann et al., 2022; Zhai et al., 2022).
+Brain decoding has yet to benefit from similar scaling.
 The challenge of cross-subject generalization has kept effective dataset sizes small, preventing the successful application of large-scale, general-purpose architectures.
-As a result, the field remains dominated by specialized models that have not yet converged to a standard, foundational approach [2, 12].
-To bridge this gap, we address the problem of subject-specific distribution shifts, which we identify as the primary obstacle to building foundation models for multi-subject problems such as EEG classification.
-We propose Subject-Conditioned Layer, an approach that decomposes the parameters of a neural network into two components: a set of base weights shared across all subjects and a subject-specific, low-rank correction that efficiently captures individual variability.
-The Subject-Conditioned Layer module is designed as a drop-in replacement for standard linear or convolutional layers, making it universally applicable to a wide range of model architectures with applications in EEG decoding and beyond.
-Our contributions are threefold: We introduce Subject-Conditioned Layer, a novel adaptive layer that explicitly separates shared knowledge from subject-specific signatures and is designed as a drop-in replacement for standard linear and convolutional layers in any neural network architecture.
-We provide qualitative evidence through embedding visualization that our layer successfully disentangles subject-invariant neural patterns from individual subject signatures, confirming the mechanism behind our method’s success.
-Through a comprehensive evaluation on the BCI Competition IV datasets, we demonstrate that both CNN and ViT like architectures equipped with Subject-Conditioned Layer outperform strong baselines: (i) a subject-agnostic model, (ii) the average of subject-specific models, and (iii) the average of subject-specific LoRA models.
-The code is publicly available  [^1] https://github.com/timonkl/SubjectConditionedLayer.
-Figure 1: Latent Representations Colored by Subject ID vs.
-Target Labels.
-The top row compares t-SNE projections of representations derived from (Left) the shared weights ($W_{\text{general}}$), (Middle) the subject-specific weights ($W_{s}$), and (Right) the fused representations from the complete Subject-Conditioned Layer.
-In this row, all points are colored by subject ID.
-The bottom row presents the same data points, but colored according to their ground-truth target labels.
-This comparison clearly shows that the sub-clusters formed by our full model (top right) correspond directly to the target classes (bottom right), demonstrating that the model learns discriminative, task-relevant features within each subject’s representation space.
-All representations are obtained from the classification token predicted by the Patched Brain Transformer, with a linear classification head used for visualization clarity.
+As a result, the field remains dominated by specialized models that have not yet converged to a standard, foundational approach (Aristimunha et al., 2025c; Guetschel et al., 2024).
+To bridge this gap, we address the problem of subject-specific distribution shifts, which we identify as the primary obstacle to building foundation models for multi-subject brain-signal applications.
+We propose the Subject-Specific Low-Rank Adapter (SuLoRA), an approach that decomposes the parameters of a neural network into two subsets: a set of base weights shared across all subjects and a subject-specific, low-rank correction that efficiently captures individual variability.
+The SuLoRA module is designed as a drop-in replacement for standard linear or convolutional layers, making it universally applicable to a wide range of model architectures with applications in EEG/MEG decoding and beyond.
+Our contributions read as follows:
+  • We introduce SuLoRA, a novel adaptive layer that explicitly separates shared knowledge from subject-specific signatures, designed as a drop-in replacement for linear and convolutional layers in any architecture.
+  • We provide qualitative evidence through embedding visualization that SuLoRA successfully disentangles subject-invariant neural patterns from individual signatures.
+  • We demonstrate through comprehensive evaluation on speech perception (MEG) and motor imagery (EEG) tasks that models equipped with SuLoRA outperform: (i) subject-agnostic models, (ii) independently trained subject-specific models, and (iii) independent subject-specific LoRA models.
+Code [^1] https://github.com/timonkl/SubjectConditionedLayer and data are publicly available.
 2 Related Work
-Our proposed method addresses the challenge of inter-subject variability in EEG decoding by learning both shared and subject-specific representations.
-This approach is closely related to established paradigms nonlinear mixed-effects (NLME) modeling and in machine learning, primarily Multi-Task Learning (MTL) and Mixture of Experts (MoE).
-Nonlinear Mixed-Effects Modeling.
-Our approach can be viewed through the lens of nonlinear mixed-effects (NLME) models [21, 22].
-In this framing, the shared network $W_{\text{general}}$ learns the fixed, population-level effects, while the subject-specific low-rank adapters $W_{s}$ capture the random effects, parsimonious deviations for each individual.
-The low-rank constraint on the adapters is analogous to the low-dimensional random effects assumption in classical NLME, which encourages partial pooling of information and borrows statistical strength across subjects.
-This perspective provides a principled justification for using LoRA [14] as a parameter-efficient method to model subject-specific variations, which is crucial when per-subject data is limited.
-Multi-Task Learning.
-In MTL, a model is trained on several related tasks simultaneously, with the goal of improving performance by leveraging shared information across tasks [5, 23, 9].
-Viewing the decoding for each subject as a distinct task, our model fits directly into the MTL framework.
-It employs a form of hard parameter sharing, where the general weight matrix, $W_{\text{general}}$, learns a common feature space shared across all subjects.
-The subject-specific weights, $W_{s}$, act as task-specific layers that adapt this shared representation to the unique neural patterns of each individual.
-This architecture allows the model to learn robust, generalizable features from the entire population while still retaining the flexibility to specialize for each subject [38].
-Mixture of Experts.
-MoE models utilize multiple expert networks and a gating mechanism that determines which expert processes a given input [15].
-Sparsely-gated MoE models, in particular, activate a subset of experts for each input, enabling the scaling of model capacity without a proportional increase in computation [31, 10].
-Our subject-aware layer can be interpreted as a specialized, non-learned MoE.
-Here, the subject identity serves as the gating signal, deterministically routing the input to its corresponding expert, the pathway defined by $W_{s}$.
-Unlike traditional MoE, our gating is not a learned function but is provided by the input via the mask M.
-This simplifies the model by removing the need to learn a routing policy and directly addresses the known source of variation (i.e., subject identity).
-Dynamical low-rank.
-Since our method explicitly models subject-specific variability as lightweight low-rank corrections, it naturally connects to the broader line of dynamical low-rank research [26, 28, 29].
-An exciting direction for future work is to explore whether subject adaptation in EEG can benefit from recent advances in geometry-aware training [36, 30], and robust compression strategies [29], thereby unifying personalization with efficient large-scale deployment.
+Scaling laws in natural language processing and computer vision have established a reliable paradigm for improving predictive performance by jointly increasing model size, training compute, and dataset scale (Kaplan et al., 2020; Hoffmann et al., 2022; Zhai et al., 2022).
+In contrast, recent work on scaling laws for neural decoding shows that performance improves with increased training data per subject, but does not scale proportionally with the number of subjects (Banville et al., 2025).
+This presents a fundamental obstacle to developing large-scale, multi-subject foundation models for brain decoding, where aggregating data across individuals is both necessary and desirable.
+Despite these challenges, Leave-One-Subject-Out (LOSO) benchmarks (Wang et al., 2020; Kwon et al., 2020; Zhang et al., 2025) demonstrate that cross-subject transfer is generally possible, although with notable performance degradation.
+Taken together, these findings suggest that brain decoding representations comprise both transferable subject-agnostic components and subject-specific components that limit performance gains when scaling across subjects.
+To address inter-subject variability in brain decoding, a wide range of approaches has been proposed.
+Early methods introduce subject-dependent bias terms or linear transformations applied to learned features (Ma et al., 2019).
+Recent studies investigate inter-subject variability via explicit adaptation mechanisms, such as subject-conditioned batch normalization and adaptive feature modulation.
+In particular, (Liang et al., 2024) learn a shared subject-invariant feature space while applying lightweight subject-specific transformations to correct individual distortions, enabling cross-subject decoding using a small amount of target-subject data.
+More recent work in speech decoding (lévy2025braintotextdecodingnoninvasiveapproach; d’Ascoli et al., 2024; Shapovalenko and Auster, 2025) introduces a subject-specific linear layer applied early in the network.
+This layer performs a channel-wise transformation that aligns subject-specific representations into a shared latent space while keeping the remaining parameters shared across subjects.
+However, this approach introduces a relatively large number of additional parameters that grow linearly with the number of subjects.
+Ablating these subject-specific layers leads to significant accuracy drops (Défossez et al., 2023b), underscoring the importance of explicitly modeling subject-dependent effects.
+Our SuLoRA directly addresses inter-subject variability by explicitly learning both shared and subject-specific representations.
+It introduces a set of global parameters augmented with lightweight, subject-specific corrections.
+This design is inspired by nonlinear mixed-effects models (Lindstrom and Bates, 1990; Pinheiro and Bates, 2000), which jointly model population-level effects and parsimonious individual-level deviations via random effects.
+The approach also naturally aligns with multi-task learning (MTL) (Caruana, 1997; Evgeniou and Pontil, 2004; Ruder, 2017), where shared structure across related tasks is leveraged while allowing task-specific variation.
+Conceptually, our approach is related to Mixture-of-Experts (MoE) models (Jiang et al., 2024a; Dai et al., 2024; Riquelme et al., 2021), which decompose computation across multiple expert networks and use a gating mechanism to select experts for each input (Jacobs et al., 1991).
+Sparsely gated MoE variants enable scaling model capacity without proportional increases in computational cost by activating only a subset of experts per example (Shazeer et al., 2017; Fedus et al., 2022).
+MoE architectures have been adapted for brain decoding: EEGMamba (Gui et al., 2024) employs a task-aware learned gating network together with a universal expert to capture shared EEG structure, while MoGE (Liu et al., 2024) and VMoGE (Ding et al., 2025) explore graph-based experts that assign channels or frequency bands to specialized experts.
+In brain decoding, subject identity provides a deterministic and known source of variation, acting as a fixed gating signal that routes each input to its corresponding expert pathway.
+Unlike conventional MoE models that rely on learned routing policies, our approach exploits this prior structure directly, simplifying the architecture while explicitly modeling the source of inter-subject variability that limits cross-subject scaling.
+Figure 1: SuLoRA architecture.
+The input $X$ is processed by a shared weight matrix $W_{\text{general}}\in\mathbb{R}^{m\times n}$ capturing subject-invariant patterns.
+Each subject s has a dedicated low-rank adapter, parameterized by matrices $A_{s}\in\mathbb{R}^{n\times r}$ and $B_{s}\in\mathbb{R}^{r\times m}$, which computes a subject-specific correction added to the shared output.
+Unlike standard MoE models, SuLoRA uses known subject identity to deterministically select the appropriate adapter at both training and inference.
+Our implementation leverages a shared full-rank expert that is always active, together with subject-specific LoRA modules.
+LoRA (Hu et al., 2021), originally introduced as a parameter-efficient fine-tuning method that adds low-rank updates to pretrained weights, has been successfully applied across modalities and tasks, offering strong empirical performance with minimal memory and compute overhead.
+Prior work has explored design choices such as rank selection (Zangrando et al., 2024; schotthöfer2024federateddynamicallowranktraining) and robust compression strategies (schotthöfer2025dynamicallowrankcompressionneural).
+In our framework, LoRA serves as the mechanism for implementing subject-specific corrections, enabling fine-grained adaptation to individual subjects while tightly controlling the number of additional parameters.
+It has also been shown that LoRA enables parameter-efficient fine-tuning of large brainwave models, achieving competitive EEG decoding performance with minimal trainable parameters (Lee et al., 2025).
 3 On the Difficulty of Multi-Subject Training
-We consider a classification problem where data is collected from multiple subjects.
+In this section, we focus solely on a classification problem in which data is collected from multiple subjects.
 Since the signal contains subject-specific characteristics, our goal is to learn a general model that captures class-discriminative structures across subjects.
@@
 This conditional independence assumption implies that $\mathcal{Z}$ serves as a sufficient statistic of $x_{sj}$ to predict $y_{sj}$.
-A key step is to verify whether this property holds for the proposed method.
 The second term of Equation (2), $p(\mathcal{Z}\mid\mathcal{X},s)$, describes how the latent representation is learned. Specifically,
@@
 4 Method
-Figure 2: Subject-Conditioned Layer.
-The input $X$ is processed by a shared general weight matrix $W_{\text{general}}$.
-For each subject, a subject-specific low-rank adapter, parameterized by matrices $A_{s}$ and $B_{s}$ of rank $r$, computes a correction that is added to the output of the general weight matrix, yielding the layer output $\bar{X}$.
-4.1 Subject-Conditioned Layer
+4.1 Subject-Specific Low-Rank Adapter
 We propose a novel subject-aware layer for multi-subject machine learning tasks as
@@
 Note that this approach may lead to a reduction in batch sizes for subject-dependent weights $W_{s}$, if the number of subjects $N$ increases.
-As a remedy, adapting the total batch size was proposed in [31].
+As a remedy, adapting the total batch size was proposed in (Shazeer et al., 2017).
 4.2 Low Rank approximation for Subject-Dependent Weights
-Low-rank adaptation (LoRA) [14] is a well-established method for fine-tuning models in few-shot settings.
-Compared to full-rank training, LoRA shows superior performance on limited data [27], while also substantially reducing the number of training parameters and the associated memory consumption.
+Low-rank adaptation (LoRA) (Hu et al., 2021) is a well-established method for fine-tuning models in few-shot settings.
+Compared to full-rank training, LoRA shows superior performance on limited data (Schotthöfer et al., 2025), while also substantially reducing the number of training parameters and the associated memory consumption.
 We define subject-specific weights as low-rank matrix decompositions
 $W_{s}^{\top}:=A_{s}B_{s},$ (5)
-with $A_{s}\in\mathbb{R}^{n\times r},\;B_{s}\in\mathbb{R}^{r\times m}$ and rank $r$. Similarly to [14], we initialize $A_{s}=\mathcal{N}(0,\sigma^{2})$, and $B_{s}=0$ for all subjects $s$.
-4.3 Regularization of Subject-Dependent Weights
+with $A_{s}\in\mathbb{R}^{n\times r},\;B_{s}\in\mathbb{R}^{r\times m}$ and rank $r$.
+Unlike standard LoRA, we initialize $A_{s}\sim\mathcal{N}\left(0,\frac{2}{r}\right)$ and $B_{s}\sim\mathcal{N}(0,0.01^{2})$ for all subjects $s$.
+Initializing $B_{s}=0$, as in standard LoRA, led to reduced performance, likely due to vanishing gradients through $A_{s}$ in early training.
+4.3 Scaling of Subject-Specific Adapters
+Similar to the original LoRA, we introduce a scaling of $W_{s}$ with $\frac{\alpha}{r}$, where $\alpha$ is a hyperparameter.
+This scaling adjusts the effective learning rate of the adapter matrices relative to the general weights.
+For $\alpha<r$, the adapters are updated conservatively, encouraging reliance on $W_{\text{general}}$, in contrast to $\alpha>r$, where the adapters are emphasized.
+4.4 Regularization of Subject-Specific Adapters
 A central challenge is to ensure that the general weight matrix $W_{\text{general}}$ learns shared patterns while the subject-specific weights $W_{s}$ only capture individual deviations.
 To prevent $W_{s}$ from learning general representations, we employ a regularization strategy.
-First, we constrain the complexity of the subject-specific corrections by enforcing a relatively small rank $r$, as defined in Equation 5.
-See Tables 2 and 3 for exact values.
-Second, we introduce an optimization-based regularization method by scaling $W_{s}$ with $\frac{\alpha}{r}$, where $\alpha$ is a hyperparameter.
-This scaling adjusts the effective learning rate of the adapter matrices relative to the general weights.
-For $\alpha<r$, the adapters are updated conservatively, encouraging reliance on $W_{\text{general}}$.
-In contrast to $\alpha>r$, where the adapters are emphasized.
+We constrain the complexity of the subject-specific corrections by enforcing a relatively small rank $r$, as defined in Equation 5.
+See Tables 4 and 5 for exact values.
 The weights $W_{s}$ could also be regularized by adding a weight regularization term (e.g., an $L_{2}$ penalty) to the loss function.
 We did not pursue this here, opting instead to demonstrate the capabilities of the LoRA only approach.
-5 Extension to Convolutional Layers
+5 Decoding Tasks and Datasets
+We evaluate our approach on two distinct neural decoding tasks: speech perception from MEG recordings and motor imagery classification from EEG data.
+5.1 Speech Perception
+We employ the MEG dataset (Gwilliams et al., 2023), a high-quality MEG dataset for natural speech processing.
+The data set consists of 208-channel MEG recordings from 27 English-speaking participants who listened to four fictional stories from the Masc corpus (Ide et al., 2010).
+The total recording duration is 56.2h.
+For each subject, data were collected in two 1-hour sessions, although the second session is missing for five subjects.
+This yields 4417 training segments and 1363 test segments with a vocabulary of 1810 and 846 unique words, respectively, with 64% word overlap between splits.
+Following (Défossez et al., 2023b), we use the same data split and preprocessing procedure, and formulate the speech-perception decoding task as a segment retrieval problem: given a short window of MEG signal, the model must identify the corresponding speech segment from a large set of candidate audio segments.
+5.2 Motor Imagery
+To complement retrieval-based speech decoding with Brain Computer Interface (BCI) classification, we additionally consider two datasets from BCI Competition IV (Tangermann et al., 2012).
+The BCI Competition IV 2a dataset (Brunner et al., 2008) consists of 22-channel EEG recordings from nine subjects.
+For each subject, data were collected in two sessions (training and evaluation), with 288 trials per session.
+The labels are balanced across four motor imagery classes: left hand, right hand, foot, and tongue.
+The BCI Competition IV 2b dataset (Leeb et al., 2007, 2008) consists of 3-channel EEG recordings from nine subjects.
+For each subject, data were collected across five sessions, three for training and two for evaluation, totaling between 680 and 760 trials.
+The labels are balanced between the classes: left-hand and right-hand.
+6 Architectures and Training Objectives
+To demonstrate the generality of SuLoRA as a drop-in architectural component, we benchmark it across three distinct modeling setups: a convolutional network, a vision transformer-based architecture, and a contrastively trained representation-learning pipeline.
+6.1 BrainMagick
+BrainMagick (Défossez et al., 2023b) addresses the challenging task of decoding speech from non-invasive MEG and EEG recordings.
+The architecture consists of two components: a Brain module and a Speech module.
+The Brain module comprises a spatial attention mechanism that encodes MEG/EEG sensors, a subject-specific linear layer, and a stack of convolutional blocks.
+The Speech module is a frozen wav2vec 2.0 speech encoder (Baevski et al., 2020).
+The Brain module is trained using a CLIP (Contrastive Language–Image Pre-Training) loss (Radford et al., 2021) to align latent brain representations with speech representations.
+A critical design choice in BrainMagick is the use of a subject-specific layer to handle substantial inter-individual variability in MEG/EEG signals.
+This layer accounts for approximately 2M of the 9.5M (for 27 subjects) trainable model parameters and scales linearly with the number of subjects.
+We replaced the subject-specific layer of the BrainMagick with SuLoRA.
+6.2 EEGNeX
+EEGNeX (Chen et al., 2024) represents the state-of-the-art among compact, convolution-only architectures for motor imagery classification, that refines the widely-adopted EEGNet (Lawhern et al., 2018).
+The network consists of five sequential blocks.
+Blocks 1 and 2 apply temporal convolutions with batch normalization to learn frequency-selective filter banks, operating analogously to learned Finite Impulse Response filters.
+Block 3 performs depthwise spatial convolution across electrode channels with max-norm constraints, followed by ELU activation, average pooling, and dropout.
+Blocks 4 and 5 introduce dilated temporal convolutions to expand the receptive field while maintaining parameter efficiency, with the final block including additional pooling and flattening operations.
+A linear layer constrained by the maximum norm serves as the classifier.
+We integrate SuLoRA by replacing standard convolutional layers while preserving depthwise spatial convolution in Block 3.
+6.3 Patched Brain Transformer
+The Patched Brain Transformer (PBT) (Klein et al., 2025a) adapts the Vision Transformer (Dosovitskiy et al., 2021) architecture for EEG decoding, that reflect recent trends in the field (Kostas et al., 2021; Jiang et al., 2024b).
+Raw EEG signals are segmented into fixed-duration patches, which are linearly projected into token embeddings.
+A learned positional embedding encodes both spatial (electrode) and temporal (time window) information, enabling the model to flexibly handle varying sensor configurations and recording durations.
+The core architecture follows the standard Transformer encoder design.
+We apply SuLoRA to the linear projections within the Transformer encoder, decomposing the query, key, value, and feed-forward weight matrices into shared and subject-specific components.
+7 Results
+We evaluate our method on diverse neural signal decoding tasks: speech perception from MEG recordings, and motor imagery classification from EEG, demonstrating the generality of our approach across different brain recording modalities and cognitive paradigms.
+Table 1: Speech decoding performance on the (Gwilliams et al., 2023) MEG dataset.
+We replace the subject-specific layer from BrainMagick (Défossez et al., 2023b) (Baseline) with SuLoRA at varying ranks $r$ and scaling factors $\alpha$.
+“BL w/o Sub.-Layer” refers to the Baseline where the subject-specific layer is replaced with the same linear layer for all subjects.
+The parameter count indicates the number of parameters in the subject-specific layer for each respective method.
+SuLoRA with $r=8$ nearly matches the Baseline performance while reducing parameters by an order of magnitude.
+SuLoRA with $r=32$ improves upon Baseline results using only half the number of parameters.
+Method subject-specific Accuracy [%] $\pm$ Std. [%]
+#Params Top 1 Top 5 Top 10
+SuLoRA r=4, $\alpha$=4 1,131k 37.56 $\pm$ 0.67 58.72 $\pm$ 0.52 67.06 $\pm$ 0.54
+r=4, $\alpha$=8 1,131k 38.85 $\pm$ 0.33 60.17 $\pm$ 0.45 68.52 $\pm$ 0.39
+r=4, $\alpha$=12 1,131k 39.04 $\pm$ 0.20 60.36 $\pm$ 0.21 68.77 $\pm$ 0.36
+r=8, $\alpha$=8 1,190k 39.82 $\pm$ 0.72 61.03 $\pm$ 0.68 69.39 $\pm$ 0.62
+r=8, $\alpha$=16 1,190k 39.61 $\pm$ 0.48 61.02 $\pm$ 0.27 69.40 $\pm$ 0.37
+r=8, $\alpha$=24 1,190k 40.32 $\pm$ 0.50 61.5 $\pm$ 0.23 69.79 $\pm$ 0.29
+r=16, $\alpha$=16 1,306k 40.50 $\pm$ 0.27 61.77 $\pm$ 0.19 69.97 $\pm$ 0.22
+r=16, $\alpha$=32 1,306k 40.50 $\pm$ 0.27 61.86 $\pm$ 0.32 70.09 $\pm$ 0.29
+r=16, $\alpha$=48 1,306k 40.68 $\pm$ 0.51 61.94 $\pm$ 0.39 70.20 $\pm$ 0.36
+[…]
+r=32, $\alpha$=64 1,540k 41.65 $\pm$ 0.76 62.63 $\pm$ 0.84 70.76 $\pm$ 0.76
+r=32, $\alpha$=96 1,540k 41.44 $\pm$ 0.32 62.67 $\pm$ 0.31 70.60 $\pm$ 0.18
+r=64, $\alpha$=64 1,006k 40.97 $\pm$ 0.44 62.15 $\pm$ 0.49 70.30 $\pm$ 0.49
+r=64, $\alpha$=128 1,006k 41.74 $\pm$ 0.49 62.90 $\pm$ 0.43 70.98 $\pm$ 0.45
+r=64, $\alpha$=192 1,006k 41.94 $\pm$ 0.92 63.20 $\pm$ 0.65 71.18 $\pm$ 0.61
+r=64, $\alpha$=256 1,006k 41.81 $\pm$ 0.21 63.00 $\pm$ 0.11 70.95 $\pm$ 0.21
+r=96, $\alpha$=192 1,473k 41.62 $\pm$ 0.13 62.73 $\pm$ 0.17 70.85 $\pm$ 0.16
+Baseline 1,968k 40.60 $\pm$ 0.18 61.84 $\pm$ 0.21 70.00 $\pm$ 0.21
+BL w/o Sub.-Layer 1,1 73k 19.23 $\pm$ 0.41 36.69 $\pm$ 0.54 45.48 $\pm$ 0.69
+Table 2: Zero-Shot.
+Prediction accuracy (%) under Leave-One-Subject-Out evaluation of EEGNeX (Chen et al., 2024) on BCI Competition IV-2a using SuLoRA, compared to the subject-agnostic baseline.
+SuLoRA achieves competitive accuracy relative to the baseline when relying solely on the general weights, indicating that general EEG features are effectively captured by these weights (highest values in bold).
+Transferred Adapters.
+Prediction accuracy (%) under Leave-One-Subject-Out evaluation when transferring an adapter trained on one training subject to predict the held-out subject.
+Adapter $i$ corresponds to subject $i$; therefore, diagonal entries are omitted.
+Transferred adapters outperform the baseline, enabling a form of few-shot learning by either selecting an appropriate adapter or using it as an initialization for LoRA fine-tuning.
+(Values shown in bold when higher than the Zero-Shot results for SuLoRA and baseline.)
+Method Leave Out Subject
+Accuracy [%] $\pm$ Std. [%]
+Zero-Shot
+S1 S2 S3 S4 S5 S6 S7 S8 S9
+[…]
+EEGNeX SuLoRA 37.04 $\pm$ 4.02 25.23 $\pm$ 0.16 56.37 $\pm$ 4.52 32.76 $\pm$ 0.16 25.12 $\pm$ 0.16 28.13 $\pm$ 1.50 28.93 $\pm$ 0.59 58.22 $\pm$ 1.99 45.14 $\pm$ 2.04
+Adapter Transferred Adapters
+S1 S2 S3 S4 S5 S6 S7 S8 S9
+1 n/a 28.01 $\pm$ 0.82 67.94 $\pm$ 3.23 34.14 $\pm$ 1.61 24.65 $\pm$ 0.28 29.28 $\pm$ 1.89 34.49 $\pm$ 1.93 46.41 $\pm$ 4.03 34.95 $\pm$ 0.87
+2 45.83 $\pm$ 3.71 n/a 28.59 $\pm$ 1.66 31.71 $\pm$ 2.05 28.70 $\pm$ 2.90 26.39 $\pm$ 0.57 16.43 $\pm$ 2.36 31.25 $\pm$ 3.97 24.19 $\pm$ 1.66
+3 61.81 $\pm$ 3.00 28.12 $\pm$ 0.85 n/a 27.43 $\pm$ 0.98 26.97 $\pm$ 1.56 31.14 $\pm$ 1.45 29.98 $\pm$ 1.56 48.84 $\pm$ 5.42 39.01 $\pm$ 3.70
+4 47.34 $\pm$ 5.57 22.92 $\pm$ 1.02 33.91 $\pm$ 5.70 n/a 26.27 $\pm$ 1.31 31.60 $\pm$ 2.32 22.80 $\pm$ 2.54 54.86 $\pm$ 2.60 34.26 $\pm$ 2.01
+[…]
+6 36.46 $\pm$ 0.29 25.46 $\pm$ 2.27 35.88 $\pm$ 6.29 37.96 $\pm$ 0.16 25.00 $\pm$ 0.29 n/a 30.21 $\pm$ 1.98 45.95 $\pm$ 2.87 43.52 $\pm$ 4.61
+7 52.43 $\pm$ 3.00 24.65 $\pm$ 2.14 45.95 $\pm$ 6.12 26.62 $\pm$ 0.87 25.35 $\pm$ 0.00 26.62 $\pm$ 1.61 n/a 32.18 $\pm$ 0.91 32.18 $\pm$ 3.99
+8 54.51 $\pm$ 2.88 29.05 $\pm$ 2.87 59.84 $\pm$ 5.67 43.06 $\pm$ 2.42 24.31 $\pm$ 0.57 28.94 $\pm$ 2.01 28.12 $\pm$ 3.07 n/a 45.14 $\pm$ 2.47
+9 29.05 $\pm$ 0.59 25.35 $\pm$ 0.28 31.25 $\pm$ 1.99 26.27 $\pm$ 1.56 25.00 $\pm$ 0.00 28.36 $\pm$ 1.61 23.26 $\pm$ 2.47 41.09 $\pm$ 4.13 n/a
+7.1 Application to Speech Decoding
+To demonstrate that SuLoRA can serve as an efficient drop-in replacement, we reproduced the speech decoding results from  (Défossez et al., 2023b) on the MEG dataset (Gwilliams et al., 2023) (see Sec.
+5.1).
+We used the publicly available BrainMagick GitHub repository (Défossez et al., 2023a) with identical hyperparameters and random seeds {2036, 2037, 2038}.
+Unlike the original work, which used gradient accumulation across two (Nvidia V100) GPUs, we trained the model on a single (Nvidia H100) GPU.
+Since contrastive losses draw negative samples only within-GPU batches, our single-GPU configuration provides access to the full batch of negatives.
+In Table 1, we compare SuLoRA with the BrainMagick model (Baseline), that achieves $40.60\%$ a top-1 accuracy (top-10: $70.00\%$) using 1,968k parameters in the subject-specific layer.
+Replacing the subject-specific layer with the same linear layer for all subjects, causes performance to collapse to top-1 accuracy of $19.23\%$ (top-10: $45.48\%$), underscoring the importance of modeling inter-subject variability in neural decoding tasks.
+SuLoRA provides a favorable accuracy-efficiency trade-off across the parameter spectrum.
+For $r=8$ and $\alpha=24$, it closely matched the baseline accuracy of $40.32\%$ top-1 (top-10: $69.79\%$) with only 190k parameters, a reduction of an order of magnitude compared to the baseline.
+Peak performance is achieved at $r=64$ with $\alpha=192$, reaching $41.94\%$ top-1, $63.20\%$ top-5, and $71.18\%$ top-10 accuracy.
+Notably, this configuration still uses roughly half the parameters of the baseline (1,006k vs.
+1,968k).
+Further increasing rank does not improve performance.
+The scaling factor $\alpha$ plays a crucial role in performance.
+For a fixed rank, we observe consistent improvements when increasing $\alpha$ up to $\alpha=3r$, after which performance plateaus or slightly decreases.
+7.2 Leave-One-Subject-Out
+Figure 2: Latent Representations Colored by Subject ID vs.
+Target Labels.
+The top row compares t-SNE projections of representations derived from (Left) the shared weights ($W_{\text{general}}$), (Middle) the subject-specific weights ($W_{s}$), and (Right) the fused representations from the complete SuLoRA.
+In the top row, all points are colored by subject ID.
+The bottom row presents the same data points, but colored according to their ground-truth target labels.
+This comparison clearly shows that the sub-clusters formed by our full model (top right) correspond directly to the target classes (bottom right), demonstrating that the model learns discriminative, task-relevant features within each subject’s representation space.
+All representations are obtained from the classification token predicted by the Patched Brain Transformer, with a linear classification head used for visualization clarity.
+Table 2 reports prediction accuracy under a Leave-One-Subject-Out (LOSO) evaluation on the BCI Competition IV-2a dataset using EEGNeX.
+In each fold, one subject is held out for testing while the model is trained on the remaining eight subjects.
+We compare a subject-agnostic baseline, SuLoRA in a zero-shot setting, and transferred adapters trained on individual source subjects.
+In the zero-shot regime, predictions using SuLoRA rely solely on the shared, subject-independent weights.
+SuLoRA achieves accuracy comparable to the subject-agnostic baseline across most held-out subjects.
+While performance varies across individuals, SuLoRA matches or exceeds the baseline for some subjects (e.g., Subjects 3 and 9), indicating that the shared representations capture meaningful and transferable EEG structure even without subject-specific adaptation.
+We further assess the transferability of subject-specific adapters by applying adapters trained on a source subject to unseen target subjects.
+Transferred adapters consistently outperform both zero-shot SuLoRA and the subject-agnostic baseline for many subject pairs.
+Notably, certain adapters yield substantial improvements for specific held-out subjects (e.g., Adapter 1 for Subject 3 and Adapter 8 for Subject 4), highlighting the presence of cross-subject structure that can be exploited without retraining the full model.
+Overall, these results demonstrate that SuLoRA supports zero-shot generalization for cross-subject adaptation, while the results of transferred adapters indicate a few-shot personalization without retraining or initialization for LoRA fine-tuning.
+Table 3: Numerical Results.
+Performance comparison across model variants: a subject-agnostic model (no personalization), subject-specific models (trained independently per subject), LoRA (independent low-rank model per subject), and SuLoRA.
+Method Total Active BCI Comp. IV2a BCI Comp. IV2b
+# Params # Params (4 Classes) (2 Classes)
+EEGNeX Subject-Agnostic (Chen et al., 2024) 155 972 55 972 55.00% $\pm$ 0.66% 74.28% $\pm$ 0.39%
+EEGNeX Subject-Specific 155 972 55 972 59.36% $\pm$ 14.42% 75.92% $\pm$ 14.32%
+EEGNex LoRA 181 124 10 980 53.60% $\pm$ 2.91% 69.64% $\pm$ 0.27%
+EEGNeX SuLoRA 134 884 64 740 64.72% $\pm$ 0.56% 76.48% $\pm$ 0.25%
+PBT Subject-Agnostic (Klein et al., 2025a) 1 867 460 867 460 50.82% $\pm$ 0.85% 75.49% $\pm$ 1.10%
+PBT Subject-Specific 1 867 460 867 460 44.16% $\pm$ 12.19% 75.63% $\pm$ 14.23%
+PBT LoRA 1 680 736 135 712 25.81% $\pm$ 0.52% 53.70% $\pm$ 0.80%
+PBT SuLoRA 1 480 612 935 588 54.51% $\pm$ 0.30% 76.36% $\pm$ 0.23%
+7.3 Representation Structure and Subject Conditioning
+Figure 2 presents t-SNE visualizations (van der Maaten and Hinton, 2008) of the learned representations to analyze the effect of subject conditioning on the latent space.
+We compare embeddings derived from the general weights $W_{\text{general}}$, the subject-specific weights $W_{s}$, and the fused representations produced by the complete SuLoRA layer.
+All embeddings are extracted from the classification token predicted by the Patched Brain Transformer, using a linear classification head for visualization.
+The top row of Figure 2 colors embeddings by subject identity, while the bottom row colors the same points by ground-truth target labels.
+Representations obtained from the general weights exhibit substantial overlap across subjects, with limited separation by either subject or class.
+Subject-specific representations show clearer grouping by subject identity, but this structure does not consistently align with the target labels.
+In contrast, the fused representations produced by the full SuLoRA display well-defined sub-clusters within each subject’s embedding space.
+These sub-clusters correspond closely to the ground-truth target labels, as evidenced by the alignment between the top-right and bottom-right panels of Figure 2.
+This indicates that the proposed model learns task-relevant, discriminative features while preserving subject-dependent structure in the representation space.
+Overall, these visualizations provide qualitative evidence that SuLoRA improves the organization of the latent space by reducing inter-subject variability while enhancing class separability within each subject.
+7.4 Application to Motor Imagery EEG Decoding
+We evaluate the SuLoRA on motor imagery EEG decoding by integrating it into two distinct architectures: EEGNeX (see Sec.
+6.2) and the Patched Brain Transformer (see Sec.
+6.3).
+The following experiment is conducted on the BCI Competition IV-2a and IV-2b datasets (see Sec.
+5.2).
+All numerical results are reported as the mean and standard deviation of three reproducible runs with random seeds $\{1,2,3\}$.
+Further details on the experimental setup are provided in Appendix B.1.
+First, to establish the necessity of subject-specific adaptation, we compare against a Subject-Agnostic Model.
+This baseline is a standard architecture trained on data collected from all subjects, without personalization.
+Our method surpasses this baseline, confirming that a one-size-fits-all approach is insufficient for high performance in EEG decoding, Table 3.
+Next, we demonstrate the benefit of knowledge sharing by comparing our method against the common practice of training fully independent Subject-Specific Models.
+This baseline trains a separate model from scratch for each individual.
+In Table 3, we report the mean performance across all subjects (the complete results are provided in Table 7).
+SuLoRA consistently outperforms the average of those independently trained models.
+This finding highlights that the shared backbone ($W_{\text{general}}$) enables effective knowledge transfer across subjects, yielding better generalization than isolated training on limited single-subject datasets.
+Finally, to validate our specific architectural design, we introduce a crucial parameter control: LoRA EEGNex.
+For each subject, a separate low-rank model is trained from scratch.
+This baseline allows us to test if the performance gain is simply an artifact of having a certain number of trainable parameters per subject.
+SuLoRA’s superiority in this comparison is a critical finding.
+It proves that our hybrid architecture is more effective than training isolated, parameter-efficient models.
+This validates that the structure of our approach is important for its performance.
+8 Conclusion
+We presented the Subject-Specific Low-Rank Adapter, a method for multi-subject neural signal decoding that explicitly separates shared and subject-specific sources of variation.
+The method integrates into existing architectures with minimal modification, requiring only replacement of standard linear or convolutional layers.
+SuLoRA addresses a fundamental obstacle to scaling foundation models for brain decoding: inter-subject distribution shift.
+The method decomposes network weights into a general component that is learned across all subjects and lightweight, low-rank subject-specific adapters.
+Our experiments across MEG speech perception and EEG motor imagery tasks demonstrate that SuLoRA consistently outperforms both subject-agnostic baselines and independently trained subject-specific models.
+In speech decoding, SuLoRA achieves superior accuracy while reducing subject-specific parameters by half compared to previous studies.
+In the motor imagery classification task, it improves performance for both CNN and transformer architectures across BCI datasets.
+The visualization of learned representations confirms that SuLoRA organizes the latent space by reducing inter-subject variability while enhancing class separability within each subject.
+More broadly, the design underlying SuLoRA may extend to other multi-source learning problems characterized by systematic distribution shifts across data sources.
+9 Limitations
+By design, SuLoRA requires subject identity at both training and inference time.
+This prevents direct application to entirely unseen subjects without adaptation.
+Although our transferred adapter experiments suggest that cross-subject structure exists, leveraging this for true zero-shot generalization to new subjects remains future work.
+The selection of the appropriate rank $r$ and the scaling factor $\alpha$ requires parameter tuning, and optimal values can vary between tasks and architectures.
+Our experiments suggest $\alpha\approx 3r$ works well, but an exact selection criterion remains an open question.
+Our evaluation focuses on datasets with 9–27 subjects.
+Although the results are consistent across different modalities and architectures, validating SuLoRA on a larger scale, such as with hundreds of subjects, would further strengthen the motivation for the foundation model.
+Finally, the total parameter count increases with the number of subjects, although modestly due to the low-rank constraint.
+For applications involving a substantial number of subjects, it may be essential to employ strategies like sharing adapters among similar subjects or utilizing hierarchical adapter structures
+Impact Statement
+This work aims to advance the fields of machine learning and neuroscience, in particular by improving the generalization of models across individuals in neural signal decoding.
+The proposed methods are intended to increase robustness and sample efficiency in learning from human brain data, which benefit research in brain–computer interfaces and cognitive neuroscience.
+The work does not introduce new data collections, and we do not anticipate negative ethical or societal consequences beyond those commonly associated with machine learning methods applied to brain signals.
+Acknowledgments
+This project has received funding from the German Federal Joint Committee (Grant 01VSF23017), from the European Regional Development Fund (grants timing Matters and IntelAlgen) under the European Union’s Horizon Europe Research and Innovation Program, and from the German Research Foundation DFG (grants GRK 2297 and 537063406), which we gratefully acknowledge.
+The work of Steffen Schotthöfer is sponsored by the Applied Mathematics Progrm at the Office of Advanced Scientific Computing Research, U.S.
+Department of Energy, and performed at the Oak Ridge National Laboratory, which is managed by UT-Battelle, LLC under Contract No.
+DE-AC05-00OR22725 with the U.S.
+Department of Energy.
+The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.
+The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
+This manuscript has been authored by UT-Battelle, LLC under Contract No.
+DE-AC05-00OR22725 with the U.S.
+Department of Energy.
+The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.
+The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan(http://energy.gov/downloads/doe-public-access-plan).
+Appendix A Extension to Convolutional Layers
 While the formulation in Equation 4 is presented for linear transformations, it can be seamlessly extended to convolutional layers.
 A convolution operation is itself a linear transformation, where the weight matrix $W$ is replaced by a convolutional kernel $W$.
-The Subject-Conditioned convolutional layer can thus be expressed as:
+The SuLoRA layer can thus be expressed as:
 $\bar{X}=\sigma\left(X*\mathcal{W}_{\text{general}}+\textstyle\sum_{s=1}^{N}M_{s}\cdot(X*\mathcal{W}_{s})\right),$ (6)
@@
 This convolutional decomposition allows for parameter-efficient adaptation for each subject while maintaining the inductive biases of convolutional layers.
-6 Data
-We conduct our experiment on two Brain Computer Interface datasets from BCI Competition IV [32].
-The data are publicly available [^2] http://bnci-horizon-2020.eu/database/data-sets with the MOABB-Package [7, 1].
-The BCI Competition IV 2a dataset [4] consists of 22-channel EEG recordings from nine subjects.
-For each subject, data were collected in two sessions (training and evaluation), with 288 trials per session.
-The labels are balanced across four motor imagery classes: left hand, right hand, foot, and tongue.
-The BCI Competition IV 2b dataset [20, 19] consists of 3-channel EEG recordings from nine subjects.
-For each subject, data were collected across five sessions, three for training and two for evaluation, totaling between 680 and 760 trials.
-The labels are balanced between the classes: left hand and right hand.
-7 Results
-We evaluate our proposed method, Subject-Conditioned Layer, by integrating it into two distinct EEG decoding architectures: EEGNeX [6], a compact Convolutional Neural Network (CNN), and the Patched Brain Transformer (PBT) [17], a Vision Transformer (ViT) based architecture [8] that reflect recent trends in the field [18, 16].
-These models demonstrate the versatility of Subject-Conditioned Layer across diverse architecture types.
-All experiments were conducted on the BCI Competition IV-2a and IV-2b datasets.
-Further details on the experimental setup are in Appendix A.
-7.1 Qualitative Analysis
-To confirm our hypothesis that our Subject-Conditioned Layer learns to decouple shared from subject-specific information, we visualize the PBT’s latent representations using t-SNE [33], see Figure 1.
-The visualizations reveal a clear separation of the learned representations.
-As shown in the figure, embeddings from the shared weights ($W_{\text{general}}$) are unstructured by subject (left), whereas those from the subject-specific weights ($W_{s}$) form distinct, individual-specific clusters (middle).
-The full model (right) integrates both aspects: it preserves the subject-level clustering while also forming meaningful sub-clusters corresponding to target labels, Figure 1, demonstrating that it retains task-relevant information.
-This demonstrates that the low-rank corrections $W_{s}$ have effectively isolated the idiosyncratic signatures of each subject.
-Together, these visualizations provide compelling evidence for our method’s proposed mechanism.
-7.2 Numerical Results
-To validate our approach, we compare Subject-Conditioned Layer against several modeling strategies, with results summarized in Table 1.
-All numerical results are reported as the mean and standard deviation of three reproducible runs with random seeds $\{1,2,3\}$.
-First, to establish the necessity of subject-specific adaptation, we compare against a Subject-Agnostic Model.
-This baseline is a standard architecture trained on data collected from all subjects, without personalization.
-Our method surpasses this baseline, confirming that a one-size-fits-all approach is insufficient for high performance in EEG decoding, Table 1.
-Next, we demonstrate the benefit of knowledge sharing by comparing our method against the common practice of training fully independent Subject-Specific Models.
-This baseline trains a separate model from scratch for each individual.
-In Table 1, we report the mean performance across all subjects (the complete results are provided in Table 4).
-Subject-Conditioned Layer consistently outperforms the average of those independently trained models.
-This finding highlights that the shared backbone ($W_{\text{general}}$) enables effective knowledge transfer across subjects, yielding better generalization than isolated training on limited single-subject datasets.
-Finally, to validate our specific architectural design, we introduce a crucial parameter control: Subject-Specific LoRA Models.
-For each subject, a separate low-rank model is trained from scratch.
-This baseline allows us to test if the performance gain is simply an artifact of having a certain number of trainable parameters per subject.
-Subject-Conditioned Layer’s superiority in this comparison is a critical finding.
-It proves that our hybrid architecture is more effective than training isolated, parameter-efficient models.
-This validates that the structure of our approach is important for its performance.
-Table 1: Numerical Results.
-Performance comparison across model variants: a subject-agnostic model (no personalization), subject-specific models (trained independently per subject), subject-specific LoRA models (entirely low-rank per subject), and Subject-Conditioned Layer.
-Method Total Active BCI Comp. IV2a BCI Comp. IV2b
-# Params # Params (4 Classes) (2 Classes)
-EEGNeX Subject-Agnostic [6] 155 972 55 972 55.00% $\pm$ 0.66% 74.28% $\pm$ 0.39%
-EEGNeX Subject-Specific 155 972 55 972 59.36% $\pm$ 14.42% 75.92% $\pm$ 14.32%
-EEGNex Subject-Specific LoRA 181 124 10 980 53.60% $\pm$ 2.91% 69.64% $\pm$ 0.27%
-EEGNeX Subject-Conditioned Layer 134 884 64 740 64.72% $\pm$ 0.56% 76.48% $\pm$ 0.25%
-PBT Subject-Agnostic [17] 1 867 460 867 460 50.82% $\pm$ 0.85% 75.49% $\pm$ 1.10%
-PBT Subject-Specific 1 867 460 867 460 44.16% $\pm$ 12.19% 75.63% $\pm$ 14.23%
-PBT Subject-Specific LoRA 1 680 736 135 712 25.81% $\pm$ 0.52% 53.70% $\pm$ 0.80%
-PBT Subject-Conditioned Layer 1 480 612 935 588 54.51% $\pm$ 0.30% 76.36% $\pm$ 0.23%
-8 Limitations & Outlook
-Our work focuses on introducing Subject-Conditioned Layer and empirically demonstrating its effectiveness on the BCI Competition IV 2a and 2b datasets.
-However, our study has some limitations, which suggest directions for future research.
-A limitation is the framework’s dependence on subject-specific data to train the personalized adapters, $W_{s}$.
-This means that for previously unseen subjects, predictions can only be made using the shared weights, $W_{\text{general}}$.
-Although we have shown that such personalization enhances performance, further analysis is required to fully assess the model’s generalization capabilities and its viability for practical deployment.
-Subject-Conditioned Layer is well-suited for a pre-training and fine-tuning paradigm.
-One could first train the shared weights $W_{\text{general}}$ across a large population.
-Subsequently, for a new subject, one could freeze the powerful shared model and efficiently fine-tune only a lightweight subject-specific adapter $W_{s}$.
-This iterative application, coupled with scaling, could lead to a Foundation model for EEG Decoding.
-To tackle the challenge of unseen subjects, future research could investigate the geometry of the learned adapter space.
-By comparing the adapters of different subjects (e.g., via their similarity), we could potentially discover subject clusters or archetypes.
-This could allow us to initialize a new subject adapter from a pre-computed group adapter, enabling faster personalization or even zero-shot prediction, thereby greatly enhancing the practical utility of our method.
-Funding Acknowledgements
-This project has received funding from the European Regional Development Fund (grants timingMatters and IntelAlgen) under the European Union’s Horizon Europe Research and Innovation Program, from the German Research Foundation DFG within GRK 2297 ‘Mathematical Complexity Reduction’, and from the German Federal Joint Committee (Grant 01VSF23017), as well as German Research Foundation DFG (Grant 537063406), which we gratefully acknowledge.
-Appendix A Experiment Design
-A.1 EEGNeX
-We used the EEGNeX implementation from the Braindecode package [11, 24].
-In this model, we replaced all standard convolutional layers with the convolutional version of our method, Subject-Conditioned Layer (see Appendix 5).
+Appendix B Numerical Experiments
+B.1 Experiment Design
+EEGNeX
+We used the EEGNeX implementation from the Braindecode package (Gramfort et al., 2013; Schirrmeister et al., 2017).
+In this model, we replaced all standard convolutional layers with the convolutional version of SuLoRA, (see Appendix A).
 We made an exception for the depthwise convolutional layer (DepthwiseConv2D) in Block 3 due to implementation details.
@@
 Patched Brain Transformer
-For the PBT, we used the official GitHub repository accompanying the publication and replaced all linear layers with our proposed method, Subject-Conditioned Layer.
+For the PBT, we used the official GitHub repository accompanying the publication and replaced all linear layers with SuLoRA.
 The only exception was for the visualization in Figure 2, where we kept the final linear layer to improve clarity.
@@
 Data augmentation Time shifts
-Appendix B Detailed Numerical Results for Subject-Specific Modeling
+BrainMagick
+For BrainMagick, we used the official GitHub repository accompanying the publication and replaced the subject-specific layer with SuLoRA.
+We used the same hyperparameters as reported in the original paper for all runs.
+Specific hyperparameters of our method, respective rank $r$ and $\alpha$ are reported in Table 1.
+Table 6: Hyperparameters BrainMagick.
+Hyperparameter Value
+Signal processing High-pass filter, downsampling
+Baseline correction 0.5 s average subtraction
+Normalization Robust scaler
+Speech representation Wav2vec 2.0 (pretrained)
+Subject embedding Yes
+Convolutional blocks Stacked CNN architecture
+Activation function GELU
+Loss function Contrastive (CLIP-style)
+Optimizer Adam
+Batch size 256
+Learning rate $3\times 10^{-4}$
+Training epochs Early stopping ($\sim$50% of test set)
+Dropout Spatial attention dropout
+Segment duration 3 seconds
+Appendix C Detailed Numerical Results for Subject-Specific Modeling
 Table 7: Subject-Specific Numerical results for EEGNeX and PBT on individual subjects.
@@
 9 75.93% $\pm$ 0.43% 83.85% $\pm$ 1.98% 55.44 % $\pm$ 0.59 % 83.02% $\pm$ 1.64%
-Appendix C Compute Resources
-The shown experiments are performed on a single NVIDIA Tesla V100 GPU using PyTorch. Code is publicly available at github.com/timonkl/SubjectConditionedLayer.
+Appendix D Software and Data
+Upon acceptance, a PyTorch implementation of SuLoRA will be released on GitHub, together with a model wrapper that allows the layer to be easily incorporated into existing architectures, including models from the Braindecode package (Aristimunha et al., 2025b).
+For the baseline model implementations used in our experiments, we utilized: the EEGNeX implementation from the Braindecode package (Aristimunha et al., 2025b), the PBT implementation from the paper’s GitHub repository (Klein et al., 2025b), and the BrainMagick implementation from the paper’s GitHub repository (Défossez et al., 2023a).
+All datasets used in this study are publicly accessible, either directly from OSF (https://osf.io/ag3kj/) or via the BCI Horizon 2020 repository (http://bnci-horizon-2020.eu/database/data-sets).
+Experiments on BCI datasets are conducted within the MOABB-Package (Chevallier et al., 2024; Aristimunha et al., 2025a).
+Appendix E Compute Resources
+The experiments for EEGNeX and PBT were performed on a single NVIDIA Tesla V100 GPU using PyTorch. The Brain Magick experiments were performed on single NVIDIA H100 GPU.